# I’m Sorry Dave: How the old world of personnel security can inform the new world of AI insider risk

### v1

@@ Abstract
Our focus in this paper is on the security risk to organisations from harmful AI insiders. However, AI also has great potential to help personnel security practitioners in defending organisations. In common with all technologies (apart from nuclear weapons) AI is dual use. In defensive mode, AI could enhance protective security in many ways – for example:
  • Strengthening pre-employment screening of people (‘vetting’) by analysing large sets of open-source data and cross-comparing multiple disparate datasets within an organisation
  • Summarising large volumes of complicated data from sources like computer event logs and access logs which may contain pointers to insider risk, but which would be indigestible for a human investigator without prior processing.
  • Enabling the continuous in-trust assurance of people and other AI systems by detecting anomalous patterns of behaviour or leading indicators of emerging insider risk and providing human security practitioners with prompts and advice about possible courses of action.
We will discuss these and other potential AI-enabled defensive capabilities in a separate paper.
@@
  • Some of the principles and methods that have been developed for tackling human insider risks may also have utility when applied to AI insiders, notwithstanding the differences in their underlying mechanisms.
  • Personnel security practitioners and AI experts should join forces to improve their mutual understanding of AI insider risks and develop better methods for countering those risks.
  • Personnel security and AI insider practitioners should make greater use of behavioural science and psychology methodologies, including conducting observational studies and experiments to identify key variables.
  • More research is needed on the nature and origins of insider risk, both in humans and AI systems. The evidence base for both is flimsy.
  • Technology producers should develop foundation models that are open and transparent, where the training data has been curated to moderate the crudeness and inaccuracy of the internet.
  • Researchers should develop effective means of evaluating an AI system’s alignment, role-dependent values, ability to follow instructions, and ability to resist subversion. Such evaluations should be kept out of the public domain to prevent the results from being gamed by bad actors.
  • Governments and regulators should encourage the widespread adoption of standards for making explicit when content has been generated by AI or decisions have been assisted by AI.
  • Methods for ensuring that LLMs do not deviate from their stated goals whilst in operation should be explored and evaluated. One way of doing this would be using multiple LLMs from different vendors to check each other’s outputs and provide feedback – a technique known as ‘LLM as a judge’. [68] This method is somewhat akin to using human colleagues as detectors of potential insider activity by their peers. It might be able to give early warning of an LLM that is starting to act in potentially harmful ways.
  • Vendors should follow best practice for the secure design and deployment of AI systems.
  • Extreme caution should be applied before deploying fully autonomous AI systems.
@@ Acknowledgements
We are grateful to Dr Daniel Martin and Kevin T. for their valuable feedback on an earlier draft of this paper.

### v2

@@ Abstract
Our focus in this paper is the security risk from AI insiders. However, AI also has great potential to help in defending organisations. AI could enhance protective security in many ways – for example: strengthening pre-employment screening of people (‘vetting’) by analysing large datasets; summarising large volumes of complicated data from sources like computer event logs and access logs which may contain pointers to insider risk; enabling the continuous in-trust assurance of people and other AI systems by detecting anomalous patterns of behaviour or indicators of emerging insider risk; and providing human security practitioners with prompts and advice about possible courses of action.
@@
  • Personnel security practitioners and AI experts should join forces to improve their mutual understanding of AI insider risks and develop better methods for countering those risks. Some of the principles and methods that have been developed for tackling human insider risks may also have utility when applied to AI insiders, notwithstanding the profound differences in their underlying mechanisms.
  • More research is needed on the nature and origins of insider risk, both in humans and AI systems. The evidence base for both is flimsy. Personnel security and AI insider practitioners should make greater use of behavioural science and psychology methodologies, including conducting observational studies and experiments to identify key variables.
  • Technology producers should develop AI systems that are open and transparent, where the training data has been curated to moderate the toxicity and inaccuracy of the internet.
  • Researchers should develop effective means of evaluating an AI system’s alignment, role-dependent values, ability to follow instructions, and ability to resist subversion.
  • Methods should be developed for ensuring that LLMs do not deviate from their stated goals whilst in operation - for example, by using multiple LLMs from different vendors to check each other’s outputs and provide feedback (a technique known as ‘LLM as a judge’ [49]). This method is akin to using human colleagues as detectors of potential insider activity by their peers.
  • Extreme caution should be applied before deploying fully autonomous AI systems [50].
@@ Acknowledgements
We are grateful to Dr Daniel Martin, Kevin T. and Ollie Whitehouse for their valuable comments on an earlier draft of this paper.